# Efficient Egocentric Human Activity Recognition

### v1

@@ 3.1 Cross-modal Self-supervised Distillation
COMODO employs two separate pretrained encoders to process video and IMU data. Given a paired video-IMU window sample $(v_{i},x_{i})$ from the dataset, we extract the features using: (1) Video encoder $f_{v}$: A frozen pretrained video encoder that maps the video $v_{i}$ to a feature representation $f_{v}(v_{i})$. The video encoder remains frozen throughout training to leverage large-scale pretrained video representations. (2) IMU encoder $f_{x}$: A pretrained but trainable time-series encoder that maps the IMU signal $x_{i}$ to a feature representation $f_{x}(x_{i})$. Unlike the video encoder, the IMU encoder is continuously updated during training to distill knowledge from the video modality. Each encoder is followed by a modality-specific projection head: (1) Video projection head $f_{v}(v_{i})$ maps $h_{v}$ to a normalized embedding $z_{v}=h_{v}(f_{v}(v_{i}))$. (2) IMU projection head $f_{x}(x_{i})$ maps $h_{x}$ to a normalized embedding $z_{x}=h_{x}(f_{x}(x_{i}))$. To enforce cross-modal distribution soft alignment, both video and IMU embedding are L2-normalized.
@@
$P_{v}(i|Q)=\frac{\exp(z_{v}\cdot z_{vi}/\tau^{v})}{\sum_{j=1}^{N}\exp(z_{v}\cdot z_{vj}/\tau^{v})},$ (1)
$P_{x}(i|Q)=\frac{\exp(z_{x}\cdot z_{vi}/\tau^{x})}{\sum_{j=1}^{N}\exp(z_{x}\cdot z_{vj}/\tau^{x})},$ (2)
[…]
$\mathcal{L}_{CE}=-\sum_{i=1}^{N}P_{v}(i|Q)\log P_{x}(i|Q)$ (3)
@@
Methods Self-supervised Ego4D EgoExo4D MMEA
Acc@1 Acc@3 Acc@5 Acc@1 Acc@3 Acc@5 Acc@1 Acc@3 Acc@5
DLinear [46] 10.60 25.27 36.74 32.20 59.91 78.16 24.01 37.73 47.18
Informer [49] 48.89 72.75 81.75 74.60 96.28 99.17 45.81 67.68 78.81
TimesNet [41] 44.33 68.02 77.79 69.20 93.44 98.04 50.30 74.24 83.99
MOMENT-small [16] 57.59 75.91 82.94 79.26 97.04 99.33 84.27 94.76 96.88
Mantis [11] 58.36 76.98 83.76 84.22 97.95 99.41 93.01 98.25 99.01
MOMENT-small [16] 39.70 64.47 75.32 68.14 93.55 98.43 83.66 95.52 97.42
Mantis [11] 47.49 71.63 81.24 76.47 96.98 99.21 90.96 98.56 99.39
IMU2CLIP [31] 54.83 76.04 84.03 79.46 97.68 99.44 91.41 98.71 99.47
COMODO (Ours) 59.13 78.79 85.65 84.92 98.28 99.59 92.48 99.01 99.77
Table 1: IMU-based Human Activity Recognition results. Our method, COMODO, achieves comparable or superior performance to fully supervised models, while consistently outperforming the previous state-of-the-art cross-modal self-supervised method, demonstrating its effectiveness in cross-modal transfer learning.
3.2 Inference
@@
Once training is complete, only the IMU encoder $f_{x}$ is used for inference, while the video encoder $f_{v}$, projector networks $h_{v},h_{x}$, and the FIFO queue $Q$ are discarded. Given an IMU window $x$, the trained IMU encoder acts as a feature extractor, producing an embedding $z_{x}$. Following prior work on unsupervised representation learning [16, 45, 12], we evaluate the quality of the learned representations by training a Support Vector Machine (SVM) with an RBF kernel on the extracted features from the training subset, and then using it to predict the classes of the test subset. In this way, our representation learning distillation framework remains unsupervised, with class labels only used to train the SVM on the extracted features.
4 Experiments
4.1 Datasets and Preprocessing
@@
To evaluate the effectiveness of COMODO, we used 3 publicly available egocentric HAR datasets: Ego4D [18], EgoExo4D [19], and MMEA [42]. The statistics of the datasets are summarized in Table 2.
@@
Table 2: Dataset statistics for Ego4D, EgoExo4D and MMEA.
MMEA provides raw sensor readings, including accelerometer and gyroscope data. Before use, we normalize the IMU data by dividing it by the provided sensitivity coefficients $R_{\text{acc}}=16384$, $R_{\text{gyro}}=16.4$ to obtain real acceleration values (in g) and gyroscope angular velocity (in ${}^{\circ}/s)$. The video data in MMEA is resized to 224 $\times$ 224 pixels, and its frame rate is maintained at 25 FPS. The IMU data are 25 Hz. Since this sampling rate is significantly lower than that of the other two datasets, it results in a relatively short sequence length, with an average of 418.57 and a median of 422. To address this, we apply left padding with an attention mask or resize the sequence length according to the requirements of different IMU student models.
@@
For Ego4D and EgoExo4D, we follow the preprocessing steps of IMU2CLIP [31]. To ensure consistency across samples, we standardize the IMU data to a uniform sampling rate of 200 Hz. The videos are resized to 224 $\times$ 224 pixels and downsampled to 10 FPS. We divide each dataset into train, test, and validation sets following a 7:2:1 ratio. Since not all Ego4D videos include IMU recordings, we first exclude those without IMU data. After this initial selection, we further exclude activity categories that contain $<10$ video samples to ensure sufficient representation per class. Similar to IMU2CLIP, we segment both video and IMU data into non-overlapping, 5-second windows, discarding the final segment if its duration is $<$ 5 seconds. Given that the IMU data in these two datasets has a higher sampling rate, each 5-second segment results in a sequence length of $5\times 200=1000$ time steps. To accommodate different IMU student models, we either subsample or resize the sequences based on their specific input requirements.
4.2 Training
Setup. In the cross-modal self-supervised distillation stage, we train COMODO for 20 epochs with a learning rate of 3e-4 and a batch size of 128. In Equation 1 and 2, the video teacher tempature is set as $\tau^{v}$ = 0.1 and the IMU student temperature is $\tau^{x}$ = 0.05. All experiments are conducted on a single NVIDIA L40S GPU.
@@ 4.2 Training
Video teacher. By default, we use the pre-trained TimeSformer-base (#params: 121.4M) [2] fine-tuned on Kinetics-400111https://huggingface.co/facebook/timesformer-base-finetuned-k400 [5] to leverage the rich knowledge from video modality. Following [7], we append an MLP projector $h_{v}$ at the end of the encoder after pooling. The hidden dimension of $h_{v}$ is 2048, the last feature dimension is 128. To ensure efficient training, all video teacher networks remain frozen throughout the distillation process. Instead of computing video features on the fly, we precompute and store all video embeddings in a pickle file, allowing direct retrieval during training to reduce computational overhead.
@@
IMU student. We choose two lightweight time-series foundation models with significantly fewer learnable parameters as the IMU student network: (1) Mantis (#params: 8.11M) [11], the current state-of-the-art classification time-series foundation model which serves as the default choice, and (2) MOMENT-small (#params: 37.9M) [16]. To generate a unified IMU representation from multiple sensor channels, we apply concat pooling, where features from each IMU channel are processed separately and then concatenated. Specifically, given an IMU input sequence $x\in\mathbb{R}^{T\times C}$ with $C$ sensor channels and $T$ time steps, we apply a shared time-series model $f_{x}$ independently to each channel: $z_{c}=f_{x}(x_{c}),\forall c\in\{1,\dots,C\}$, where $x_{c}\in\mathbb{R}^{T}$ represents the input sequence for the $c$-th IMU channel, and $z_{c}\in\mathbb{R}^{D}$ is the corresponding feature representation where $D$ denotes the dimension. The final IMU embedding $z_{x}$ is obtained by concatenating all per-channel embeddings: $z_{x}=\{z_{1},z_{2},\dots,z_{C}\}\in\mathbb{R}^{C\times D}$. This approach ensures that the model captures independent channel-wise representations while maintaining structural consistency in the concatenated feature space. In Section 4.4.3, we present an ablation study showcasing results with different input pooling methods. Similar to the video teacher network, we add one additional MLP projector $h_{x}$.
4.3 Evaluation
To show the effectiveness of our proposed framework, we evaluate the representation quality of the IMU student encoder by comparing its performance against state-of-the-art time-series models and self-supervised learning methods on two downstream tasks. We use Accuracy @1, @3, and @5 to assess learned representations, as these metrics capture both the top-ranked prediction accuracy and the overall ranking quality within the predicted similarity distribution.
4.3.1 IMU-based Human Activity Recognition
@@
This task aims to classify human activities from IMU sensor data for efficient and reliable recognition on resource-constrained wearables. Effective IMU-based HAR models should capture motion semantics and generalize robustly across diverse real-world sensor readings. We benchmark the classification on three egocentric human activity recognition datasets. Table 1 categorizes methods into supervised and Self-supervised settings:
@@
Supervised methods. We include state-of-the-art time-series models as our supervised baselines: DLinear [46], Informer [49], TimesNet [41], MOMENT [16], and Mantis [11]. These models are fine-tuned on labeled IMU data.
Self-supervised methods. We evaluate pretrained time-series foundation models as our self-supervised baselines: MOMENT-small [16] and Mantis [11], which are assessed in a zero-shot manner without fine-tuning or distillation on the target IMU dataset. Additionally, we include IMU2CLIP [31] as our contrastive learning baseline, which leverages cross-modal self-supervised knowledge transfer to allow the IMU student model to acquire knowledge from the video modality, similar to COMODO. For these self-supervised methods, we follow the evaluation protocol in Section 3.2. To ensure a fair comparison, we re-implemented IMU2CLIP, preserving its core contrastive learning framework, and replaced the original CLIP-based teacher with our video teacher network and substituted IMU2CLIP’s student encoder with our IMU student network.
Results. As shown in Table 1, COMODO achieves the highest accuracy on most metrics across all three datasets. On Ego4D, it outperforms the state-of-the-art cross-modal self-supervised baseline IMU2CLIP by 4.3% in Acc@1 and matches the best supervised fine-tuned time-series baseline Mantis with a 0.77% improvement. On EgoExo4D, it outperforms IMU2CLIP by 5.46% in Acc@1 and exceeds supervised fine-tuned Mantis by 0.7%. On MMEA, COMODO surpasses IMU2CLIP by 1.07% in Acc@1 while closely matching the best supervised method, falling short by only 0.53%. Notably, COMODO achieves the highest scores in Acc@3 and Acc@5, indicating that it captures richer semantic representations, leading to a better overall prediction distribution. These results highlight the effectiveness of COMODO in leveraging cross-modal knowledge transfer to enhance IMU-based HAR. Furthermore, its ability to match or even surpass fully supervised fine-tuned models demonstrates its potential as a scalable and label-efficient solution for HAR and other related tasks.
Train Test Method Acc@1 Acc@3 Acc@5
EgoExo4D Ego4D Mantis 52.58 75.55 83.83

### v2

@@ 3.1. Cross-modal Self-supervised Distillation
COMODO employs two separate pretrained encoders to process video and IMU data. Given a paired video-IMU window sample $(v_{i},x_{i})$ from the dataset, we extract the features using: (1) Video encoder $f_{v}$: A frozen pretrained video encoder that maps the video $v_{i}$ to a feature representation $f_{v}(v_{i})$. The video encoder remains frozen throughout training to leverage large-scale pretrained video representations. (2) IMU encoder $f_{x}$: A pretrained but trainable time-series encoder that maps the IMU signal $x_{i}$ to a feature representation $f_{x}(x_{i})$. Unlike the video encoder, the IMU encoder is continuously updated during training to distill knowledge from the video modality. To address the modality heterogeneity and potential differences in output dimensionality, we append a projection head after each encoder to map features into a shared embedding space: (1) The video projection head $h_{v}$ maps $f_{v}(v_{i})$ to a normalized embedding $z_{v}=h_{v}(f_{v}(v_{i}))$. (2) The IMU projection head $h_{x}$ maps $f_{x}(x_{i})$ to a normalized embedding $z_{x}=h_{x}(f_{x}(x_{i}))$. To enforce cross-modal distribution soft alignment, both video and IMU embedding are L2-normalized.
@@
(1) $P_{v}(i|Q)=\frac{\exp(z_{v}\cdot z_{vi}/\tau^{v})}{\sum_{j=1}^{N}\exp(z_{v}\cdot z_{vj}/\tau^{v})},$
(2) $P_{x}(i|Q)=\frac{\exp(z_{x}\cdot z_{vi}/\tau^{x})}{\sum_{j=1}^{N}\exp(z_{x}\cdot z_{vj}/\tau^{x})},$
where $P_{v}(i|Q)$ and $P_{x}(i|Q)$ represent the similarity score distributions of the video and IMU embeddings, respectively. $\tau^{v}$ and $\tau^{x}$ are temperature scaling factors for the video and IMU, and $\left(\cdot\right)$ represents the inner product between the two embeddings. To encourage the IMU encoder to capture semantic relationships present in the video modality, we align the similarity distributions of the two modalities by minimizing cross-entropy:
(3) $\mathcal{L}_{CE}=-\sum_{i=1}^{N}P_{v}(i|Q)\log P_{x}(i|Q)$
@@
3.1.1. Relation to InfoNCE and theoretical motivation.
Given a FIFO queue $Q=\{z_{v1},\ldots,z_{vN}\}$ of frozen video embeddings, we denote by $P_{v}(i)$ and $P_{x}(i)$ the similarity distributions defined in Eq. 1 and Eq. 2, respectively (with the same queue $Q$). Our objective in Eq. 3 can be rewritten as
(4) $\mathcal{L}_{CE}=-\sum_{i=1}^{N}P_{v}(i)\log P_{x}(i)\;=\;H(P_{v})+\mathrm{KL}(P_{v}\|P_{x}),$
where $H(P_{v})$ is the entropy of the video teacher distribution and is constant with respect to the parameters of $f_{x}$ (and $h_{x}$) since $f_{v}$ is frozen. Therefore, COMODO performs similarity distribution alignment by minimizing $\mathrm{KL}(P_{v}\|P_{x})$, i.e., distilling the structure of video teacher’s similarity space rather than enforcing only instance-wise alignment.
InfoNCE as a special case. As discussed above, the most recent entry in the FIFO queue corresponds to the current teacher embedding (i.e., $z_{vN}=z_{v}$), yielding the maximal similarity before softmax. When $\tau^{v}\rightarrow 0$, the teacher similarity distribution $P_{v}(i)$ becomes increasingly peaked and smoothly approaches a one-hot target with $P_{v}(N)\rightarrow 1$ and $P_{v}(i\neq N)\rightarrow 0$. In this extreme case, $\mathcal{L}_{CE}$ reduces to the “hard” contrastive objective
(5) $\mathcal{L}_{NCE}\;=\;-\log\frac{\exp(z_{x}\cdot z_{v}/\tau^{x})}{\sum_{j=1}^{N}\exp(z_{x}\cdot z_{vj}/\tau^{x})},$
which is the standard InfoNCE-style cross-modal contrastive learning.
Why the soft distribution helps in video-to-IMU distillation. The gradient of $\mathcal{L}_{CE}$ with respect to the IMU embedding $z_{x}$ is
(6) $\frac{\partial\mathcal{L}_{CE}}{\partial z_{x}}=\frac{1}{\tau^{x}}\left(\sum_{i=1}^{N}P_{x}(i)\,z_{vi}-\sum_{i=1}^{N}P_{v}(i)\,z_{vi}\right),$
which encourages $z_{x}$ to match the video teacher’s similarity distribution over the queue, thereby preserving the structure of the video teacher’s similarity space. In contrast, InfoNCE corresponds to replacing $\sum_{i=1}^{N}P_{v}(i)\,z_{vi}$ by the single positive $z_{v}$, (i.e., it treats all non-positives as equally dissimilar and pushes them away uniformly). For heterogeneous modalities (video vs. IMU), many “negatives” can be semantically close under the video teacher, (i.e., $P_{v}(i)>0$ for several $i\neq N$). In this case, the “hard” InfoNCE target (one-hot on the positive) enforces a gradient that ignores these teacher-indicated neighbors, whereas COMODO provides a soft, distribution-level target and updates $z_{x}$ to fit the teacher’s similarity distribution via $\sum_{i=1}^{N}P_{v}(i)\,z_{vi}$. Consequently, COMODO provides a distribution-level supervisory signal that preserves the structure of the video teacher’s similarity space rather than collapsing it into a single positive instance, which is well aligned with cross-modal distillation under semantic overlap among queue entries. In our experiments (Section 4.6.4), we empirically verify this advantage by comparing COMODO against InfoNCE baseline and L2 alignment.
3.2. Inference
@@
Once training is complete, only the IMU encoder $f_{x}$ is used for inference, while the video encoder $f_{v}$, projector networks $h_{v},h_{x}$, and the FIFO queue $Q$ are discarded. Given an IMU window $x$, the trained IMU encoder acts as a feature extractor, producing an embedding $z_{x}$. Following prior work on unsupervised representation learning (Goswami et al., 2024; Yue et al., 2022; Franceschi et al., 2019), we evaluate the quality of the learned representations by training a Support Vector Machine (SVM) with an RBF kernel on the extracted features from the training subset, and then using it to predict the classes of the test subset. In this way, our representation learning distillation framework remains unsupervised, with class labels only used to train the SVM on the extracted features.
4. Experiments
4.1. Datasets
@@
To evaluate the effectiveness of COMODO, we used 3 publicly available egocentric HAR datasets: Ego4D  (Grauman et al., 2022), EgoExo4D  (Grauman et al., 2024), and MMEA  (Xu et al., 2024), each with distinct sensor configurations and sampling rates. To ensure consistency and comparability, we apply dataset-specific preprocessing strategies, which we detail below. The statistics of the datasets are summarized in Table 2.
4.1.1. Dataset Activity Labels
We list below the activity categories available in each dataset.
Ego4D
The Ego4D dataset contains 31 activity categories after filtering classes with fewer than ten samples: cleaning / laundry, crafting/knitting/sewing/drawing/painting, cooking, ”jobs related to construction/renovation company (director of work, tiler, plumber, electrician, handyman, etc)”, household management - caring for kids, farmer, carpenter, scooter mechanic, reading books, walking on street, bike mechanic, playing games / video games, working out at home, working at desk, gardening, playing board games, cycling / jogging, watching tv, practicing a musical instrument, baker, playing with pets, eating, ”car - commuting, road trip”, potting plants (indoor), playing cards, fixing something in the home, car/scooter washing, bike, working out outside, biology experiments, indoor navigation (walking).
EgoExo4D
EgoExo4D provides 8 coarse activity categories: Rock Climbing, Cooking, Health, Bike Repair, Basketball, Music, Dance, Soccer.
MMEA
The MMEA dataset includes 32 fine-grained daily activities: upstairs, downstairs, drinking, fall, reading, sweep-floor, cut-fruits, mop-floor, writing, wipe-table, wash-hand, standing, play-phone, type-PC, eating, cooking, pick-up-phone, drop-trash, fold-colthes, walking, play-card, brush-teeth, wash-dish, moving-sth, type-phone, chat, open-close-door, ride-bike, sit-stand, take-drop-sth, shopping, watch-TV.
@@ 4.1.1. Dataset Activity Labels
Table 2. Ego4D / EgoExo4D / MMEA Statistics.
4.1.2. Preprocessing
@@
Ego4D and EgoExo4D are preprocessed following the same strategy as IMU2CLIP (Moon et al., 2023) to ensure consistency. First, we exclude video samples that do not contain IMU recordings. Then, to ensure sufficient class representation, we further remove activity categories with fewer than 10 samples. We divide the remaining data into train, test and validation sets using a 7:2:1 ratio, we also provide the split files in our code repository. The IMU signals are standardized to a uniform sampling rate of 200 Hz, and video frames are resized to 224 $\times$ 224 pixels and downsampled to 10 FPS. Both the video and IMU data are segmented into non-overlapping 5-second windows. The final segment is discarded if its duration is less than 5 seconds to ensure uniform window length. With a sampling rate of 200 Hz, each IMU segment corresponds to a fixed sequence length of 5 $\times$ 200 = 1000 time steps. To adapt to different IMU student models, we use model-specific sequence adjustments: MOMENT-small automatically subsamples sequences that exceed its input length limit, while Mantis applies its resize function to reshape all sequences to a fixed length of 512. These preprocessing steps ensure compatibility with diverse IMU models while preserving the temporal and structural integrity of the original sensor data.
MMEA provides raw sensor readings, including accelerometer and gyroscope data, sampled at 25 Hz. This relatively low sampling frequency leads to shorter IMU sequences, with an average sequence length of 418.57 and a median of 422, as visualized in Figure 4. The variance in sequence lengths within MMEA also allow us to evaluate the model’s adaptability and robustness to variable-length inputs. To convert raw IMU readings into real-world values, we normalize the data by dividing the accelerometer readings by $R_{\text{acc}}=16384$ and the gyroscope readings by $R_{\text{gyro}}$ = 16.4, resulting in real acceleration values (in g) and angular velocities (in $^{\circ}/s$). The video data in MMEA is resized to 224 $\times$ 224 pixels, and the frame rate is kept at 25 FPS. We follow the official train-validation-test splits provided by the MMEA dataset.
Figure 4. The visualization of distribution of sequence lengths in MMEA dataset.
Methods Self-supervised Ego4D EgoExo4D MMEA
A@1 A@3 A@5 A@1 A@3 A@5 A@1 A@3 A@5
DLinear 10.60 25.27 36.74 32.20 59.91 78.16 24.01 37.73 47.18
Informer 48.89 72.75 81.75 74.60 96.28 99.17 45.81 67.68 78.81
TimesNet 44.33 68.02 77.79 69.20 93.44 98.04 50.30 74.24 83.99
DeepConvLSTM 52.85 75.70 83.49 76.51 96.89 99.37 52.58 77.25 86.44
Attend 58.12 78.08 84.56 81.45 98.20 99.66 58.66 79.87 87.47
IMUGPT 2.0 (V) 4.25 13.44 18.77 44.76 56.80 73.93 7.15 18.48 27.77
IMUGPT 2.0 (V+R) 46.67 71.07 80.45 74.93 96.45 99.13 17.19 35.38 47.78
CrossHAR 50.96 72.66 81.53 77.14 96.95 99.31 84.97 96.55 97.74
MOMENT-small 57.59 75.91 82.94 79.26 97.04 99.33 84.27 94.76 96.88
Mantis 58.36 76.98 83.76 84.22 97.95 99.41 93.01 98.25 99.01
CrossHAR 30.18 56.10 70.00 67.80 94.67 98.66 84.80 96.88 98.33
MOMENT-small 39.70 64.47 75.32 68.14 93.55 98.43 83.66 95.52 97.42
Mantis 47.49 71.63 81.24 76.47 96.98 99.21 90.96 98.56 99.39
IMU2CLIP 54.83 76.04 84.03 79.46 97.68 99.44 91.41 98.71 99.47
COMODO (Ours) 59.13 78.79 85.65 84.92 98.28 99.59 92.48 99.01 99.77
Table 3. IMU-based Human Activity Recognition results: We compare COMODO against 14 baselines from both fully supervised and self-supervised approaches. Our method, COMODO, achieves comparable or superior performance to fully supervised models, while consistently outperforming the previous state-of-the-art self-supervised methods, demonstrating its effectiveness in cross-modal transfer learning. The best are indicated in bold, and the second-best are underlined. (V) denotes training with virtual-only IMU data; (V+R) denotes training with virtual + real IMU data.
4.2. Experimental Settings
Setup.
To evaluate the effectiveness of our proposed framework, we assess the representation quality of the IMU student encoder on two downstream tasks, comparing it against state-of-the-art time-series models, human activity recognition models, and self-supervised learning methods. Detailed configurations for all baselines, including sequence length and reimplementation specifics, are provided in Appendix A. We report Accuracy @1, @3, and @5, which reflect both top-1 prediction correctness and the quality of the predicted similarity distribution. In the cross-modal self-supervised distillation stage, we train COMODO for 20 epochs using a batch size of 128 and a learning rate of 3e-4. The temperature hyperparameters in Equations 1 and 2 are set to $\tau^{v}=0.1$ for the video teacher and $\tau^{x}=0.05$ for the IMU student. All experiments are conducted on a single NVIDIA L40S GPU. We fix the random seed to 42 for all experiments to ensure reproducibility.
Video teacher.
@@ 4.2. Experimental Settings
By default, we use the pre-trained TimeSformer-base (#params: 121.4M)  (Bertasius et al., 2021) fine-tuned on Kinetics-400 (Carreira and Zisserman, 2017) to leverage the rich knowledge from video modality. We append an MLP projector $h_{v}$ at the end of the encoder. The hidden dimension of $h_{v}$ is 2048, the output dimension is 128. To ensure efficient training, all video teacher networks remain frozen during distillation. Instead of computing video features on the fly, we precompute and store all video embeddings in a pickle file, allowing direct retrieval during training to reduce computational overhead.
IMU student.
@@
We choose two lightweight time-series foundation models with significantly fewer learnable parameters as the IMU student network: (1) Mantis (#params: 8.1M)  (Feofanov et al., 2025), the current state-of-the-art classification time-series foundation model which serves as the default choice, and (2) MOMENT-small (#params: 37.9M)  (Goswami et al., 2024). To generate a unified IMU representation from multiple sensor channels, we apply concat pooling, where features from each IMU channel are processed separately and then concatenated. Specifically, given an IMU input sequence $x\in\mathbb{R}^{T\times C}$ with $C$ sensor channels and $T$ time steps, we apply a shared time-series model $f_{x}$ independently to each channel: $z_{c}=f_{x}(x_{c}),\forall c\in\{1,\dots,C\}$, where $x_{c}\in\mathbb{R}^{T}$ represents the input sequence for the $c$-th IMU channel, and $z_{c}\in\mathbb{R}^{D}$ is the corresponding feature representation where $D$ denotes the dimension. The final IMU embedding $z_{x}$ is obtained by concatenating all per-channel embeddings: $z_{x}=\{z_{1},z_{2},\dots,z_{C}\}\in\mathbb{R}^{C\times D}$. This approach ensures that the model captures independent channel-wise representations while maintaining structural consistency in the concatenated feature space. In Section 4.6.3, we present an ablation study showcasing results with different input pooling methods. Similar to the video teacher network, we add one additional MLP projector $h_{x}$.
4.3. IMU-based Human Activity Recognition
@@
This task aims to classify human activities from IMU sensor data for efficient and reliable recognition on resource-constrained wearables. We benchmark the classification on three egocentric human activity recognition datasets. Table 3 categorizes methods into supervised and self-supervised settings (Table 3):
Supervised methods.
@@
We include methods for time-series and HAR as our supervised baselines: DLinear  (Zeng et al., 2023), Informer  (Zhou et al., 2021), TimesNet  (Wu et al., 2023), DeepConvLSTM (Ordóñez and Roggen, 2016), Attend and Discriminate (Abedin et al., 2021), IMUGPT 2.0 (Leng et al., 2024), CrossHAR (Hong et al., 2024), MOMENT-small  (Goswami et al., 2024), and Mantis  (Feofanov et al., 2025). These models are fine-tuned on labeled IMU data in this task.
Figure 5. Qualitative t-SNE visualization.
Self-supervised methods.
We include two pretrained time-series foundation models, MOMENT-small (Goswami et al., 2024) and Mantis (Feofanov et al., 2025). We also compare against CrossHAR (Hong et al., 2024), a hierarchical self-supervised pretraining framework for IMU-based HAR, and IMU2CLIP (Moon et al., 2023), a contrastive cross-modal baseline.
Results.
As shown in Table 3, COMODO achieves state-of-the-art performance across all three datasets, ranking first on most metrics. On Ego4D, it outperforms IMU2CLIP by 4.3% in A@1 and surpasses strong supervised models, Mantis and Attend and Discriminate, by 0.77% and 1.01%, respectively. COMODO also exceeds the performance of CrossHAR and IMUGPT 2.0 (Virtual+Real) by 8.17% and 12.46% in A@1. On EgoExo4D, COMODO leads in A@1 and A@3, outperforming IMU2CLIP by 5.46% and Mantis by 0.7%, while achieving second-best in A@5, just narrowly behind Attend and Discriminate. In this dataset, COMODO consistently maintains a significant lead over CrossHAR and both variants of IMUGPT 2.0. On MMEA, it exceeds IMU2CLIP by 1.07% in A@1, ranks second to supervised Mantis by only 0.53%, and achieves the best A@3 and A@5 scores. COMODO outperforms the finetuned CrossHAR by 7.51% and demonstrates a substantial accuracy gain over IMUGPT 2.0. Notably, COMODO consistently excels in A@3 and A@5 across all datasets, indicating its ability to capture richer semantic representations for a more accurate prediction distribution. These results validate the effectiveness of COMODO’s cross-modal knowledge transfer, demonstrating its capacity to not only advance self-supervised learning but also match or surpass fully supervised models, underscoring its potential as a scalable, label-efficient solution for HAR. To qualitatively assess the learned representations, Figure 5 shows t-SNE (Maaten and Hinton, 2008) visualizations on Ego4D and EgoExo4D. COMODO produces denser, more compact intra-class clusters with clearer inter-class boundaries, while IMU2CLIP exhibits greater overlap and dispersion.
Train Test Method A@1 A@3 A@5
EgoExo4D Ego4D CrossHAR 30.55 55.95 69.70
Mantis 52.58 75.55 83.83